OpenStudy (anonymous):

given that y varies directly with x, write the linear function relating the two variables if y=6 and x=16

OpenStudy (anonymous):

direct variation means y = kx, so 6 = k*16, which means k = 6/16 or 3/8. Therefore, the linear function is y = (3/8)*x

OpenStudy (anonymous):

Ask a question if any of that does not make sense.

OpenStudy (jhannybean):

Direct variation \[\large y=kx\]\[\large y = 6 \ , \ x= 16\]Solve for k. Plug in your values \[\large 6 = 16k\]\[\large k=\frac{6}{16} =\frac{3}{8} \]\[\large y=\frac{3}{8}x\]

OpenStudy (anonymous):

awesome thank you so much! That was the solution that I came up with, but since it really wasnt an even number I came up with two different possibilities.

OpenStudy (jhannybean):

When you're looking for a formula of direct variation, you're looking for values of x when plugged into the equation, how they would DIRECTLY effect the outcome of y. Something like that.